OpenStudy (anonymous):

logp/m=8 and logm/n=5 what is the relationship between M and P

OpenStudy (anonymous):

\[m = p \times e^{-8}\]

OpenStudy (loser66):

the question should be " what is the relationship between N and P, right?

OpenStudy (loser66):

because if it asks between M and P , \(log(\frac{p}{m}) =8\) shows it.

OpenStudy (agent0smith):

Assuming you meant N and P, Use the log law \[\Large \log_a b = x \text{ then } a^x = b\]so\[\large \log_{10} \frac{ p }{m}=8\]then\[\Large 10^8 = \frac{ p }{m }\]solving that for m gives \[\Large m = \frac{ p }{ 10^8 } \] Now try and follow similar steps for the logm/n = 5, and subsitute in the